OpenStudy (anonymous):

A line passes through the point(0,-3/2) and has a slope of 1/2 . What is the equation of the line?

OpenStudy (anonymous):

y-y1=m(x-x1) m is the slope which is 1/2 (x1,y1) is the point which is (0,-3/2)

OpenStudy (anonymous):

y-(-3/2)=1/2(x-0) y+3/2=1/2x adding -3/2 for both side y+3/2-3/2=1/2x-3/2 y=1/2x-3/2